# DRAGON: LLM-Driven Decomposition and Reconstruction Agents for Large-Scale Combinatorial Optimization

### v1

@@ 4.1.4. Feasibility check
Feasibility checking for solutions obtained by any method is straightforward. We implement a checker based on the inherent constraints of each COP type. If a solution is found to be infeasible, it is stored in the agent’s memory along with comments explaining the reason for infeasibility. This feedback helps the agent avoid repeating the same mistakes, thereby reducing search time and accelerating convergence to feasible solutions. In the results presented later, we denote an infeasible solution as “infe” and use “inf” to indicate cases where the solving time exceeds the given time limit $T_{\max}$.
@@ 4.2. Performance Evaluation
To evaluate performance on large-scale routing problems (TSP and CVRP), Table 1 presents the average optimality gap (%) across different instance size groups, comparing DRAGON with existing LLM-based solvers, including prompt-based methods (OPRO Yang et al. (2024), SGE Iklassov et al. (2024)) and code-generation approaches (LMEA Liu et al. (2024a), ReEvo Ye et al. (2025), with ReEvo(a) using Ant Colony Optimization (ACO) and ReEvo(c) employing a constructive heuristic). To ensure a fair comparison, all LLM-based methods use gpt-4o, with token usage and inference time recorded. Instances are grouped by node size for conciseness, and detailed results are listed in Supplementary Materials.
@@
As shown in Figure 4, results on TSPLIB reveal that SGE is excluded due to infeasible solutions, while OPRO is omitted as “inf” that no solution found due to its slow inference time across all groups. LMEA is able to solve up to medium-sized instances but exhibits relatively large optimality gaps. ReEvo(c) achieves strong results up to the medium-size group, whereas ReEvo(a) excels on smaller instances. Although DRAGON does not outperform across every group, it demonstrates clear superiority on larger-scale problems, benefiting from its task allocation mechanism, where decomposition and reconstruction collaborate effectively.
Figure 5. Optimality Gap (%) for different methods on CVRPLIB.
@@
These findings highlight a major limitation of pure prompt-based methods: while capable of solving small to medium-scale COPs, their performance degrade significantly as problem size increases. This is mainly due to long context lengths and high decoding cost—making token-by-token generation is unstable when encounter large input loads. Code-generation methods like ReEvo improve scalability. We tested ReEvo’s released heuristic code, including both constructive (ReEvo(c)) and ACO-enhanced (ReEvo(a)) variants. On TSPLIB, DRAGON consistently outperforms ReEvo(c) across all size groups and also surpasses ReEvo(a) on all groups beyond 500, and achieving an average gap reduction of 0.86%. On CVRPLIB, where OR-Tools fails to find feasible solutions within time limit, DRAGON leads around 3.98% lower gaps than ReEvo(a) on all cases. Overall, DRAGON demonstrates the best average performance across routing benchmarks, validating its effectiveness and scalability.
@@ 4.3.2. Impact of LLM backbones
Figure 7. Comparison of different LLM models by: (a) optimality gap, (b) number of API calls, (c) input tokens, (d) output tokens, (e) running time. We evaluate general-purpose models (gpt-4o, gpt-4.1) and reasoning-oriented models (o3, r1). Gray bars indicate incomplete results due to token limit violations on the largest instance.
@@ 4.4.1. Bin Packing Problem
Compared to routing, packing problems are generally easier for LLMs to reason about due to their more intuitive and straightforward constraints. For example, merging static segments ($\mathbf{s}$) into packed items requires less complex reasoning, allowing DRAGON to scale more effectively to larger instances. Results for EoH and EoH expert are reported directly from Liu et al. (2024b), where running time was not provided. While DRAGON achieves the lowest optimality gap, it is slower than code-generation based methods. However, code-generation approaches come with high API costs and substantial time and resources to evolve or fine-tune high-quality heuristic functions. These overheads are not captured in current comparison. Each paradigm has its strengths. For large batches of similar instances, code-generation methods are effective as the learned heuristic can be reused across the dataset. In contrast, prompt-based approaches like DRAGON are better suited for dynamic environments where instance structures vary, offering flexibility without the need for retraining or code evolution.
@@ 5. Conclusion
This paper presents DRAGON, a novel framework that leverages LLM agents to solve large-scale COPs via an iterative decomposition reconstruction process inspired by divide-and-conquer. Experiments across diverse COP domains (e.g., routing and packing) show that DRAGON consistently achieves strong performance, especially on extreme large instances where traditional solvers or basic prompt-based methods often fail due to scalability or timeouts. By decomposing complex problems into context manageable subproblems, DRAGON enables LLMs to reason effectively over intricate structures, identifying promising regions, and reconstructing feasible global solutions. This training-free, modular, and solver-agnostic approach supports flexible constraint handling and adapts easily to diverse COPs without specialized algorithm design. Although not optimized for speed, DRAGON avoids expensive code search and offers a strong balance between solution quality, adaptability, and cost, making it particularly well-suited for dynamic and large-scale settings. Future work will focus on enhancing decomposition strategies, improving prompt robustness, and integrating hybrid solvers to explore promising regions more efficiently.
Overall, DRAGON establishes a solid foundation toward general-purpose, interpretable, and scalable LLM-based optimization.
@@
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-017, AISG3-RP-2025-036-USNSF).
@@ B.1. TSPLIB Results
We then present full results on benchmarking instances from TSPLIB in Tables A3A7 add more commentary analysis
OPRO fails to return valid solutions for all groups, as some hard instances in groups exceed the runtime limit, resulting in an infi- nite gap inf. SGE consistently produces infeasible solutions (infe), likely because it lacks explicit feasibility checks, since its origi- nal design is for small problems (under 30 nodes). LMEA handles slightly larger instances but exhibits high gaps and fails beyond 2k-node due to prompt length limits and generation timeouts.

### v2

@@ 4.1.4. Feasibility check
We implement a checker based on the inherent constraints of each COP type. If a solution is found to be infeasible, it is stored in the agent’s memory along with comments explaining the reason for infeasibility. This feedback helps the agent avoid repeating the same mistakes, thereby reducing search time and accelerating convergence to feasible solutions. In the results presented later, we denote an infeasible solution as “infe” and use “inf” to indicate cases where the solving time exceeds the given time limit $T_{\max}$.
@@ 4.2. Performance Evaluation
To evaluate performance on large-scale routing problems (TSP and CVRP), Table 1 presents the average optimality gap (%) across different instance size groups, comparing DRAGON with existing LLM-based solvers, including prompt-based methods (OPRO yang2024large, SGE iklassov2024self) and code-generation approaches (LMEA liu2024large, ReEvo ye2024reevo, with ReEvo(a) using Ant Colony Optimization (ACO) and ReEvo(c) employing a constructive heuristic). To ensure a fair comparison, all LLM-based methods use gpt-4o, with token usage and inference time recorded. Instances are grouped by node size for conciseness, and detailed results are listed in Supplementary Materials.
@@
As shown in Figure 4, the results in TSPLIB reveal that SGE is excluded due to infeasibility, while OPRO is omitted as “inf” so that no solution was found due to its slow inference time in all groups. LMEA is able to solve up to medium-sized instances but exhibits relatively large optimality gaps. ReEvo(c) achieves strong results up to the medium-size group, whereas ReEvo(a) excels on smaller instances. Although DRAGON is not the best on every group, it shows clear advantages on larger-scale problems, driven by its task‑allocation mechanism, where decomposition and reconstruction work effectively together.
Figure 5. Optimality Gap (%) for methods on CVRPLIB.
@@
These findings highlight a major limitation of pure prompt-based methods: their performance degrades significantly as problem size increases. This is mainly due to long context lengths and high decoding cost—making token-by-token generation is unstable when encounter large input loads. Code-generation methods like ReEvo improve scalability. We tested ReEvo’s released heuristic code, including both constructive (ReEvo(c)) and ACO-enhanced (ReEvo(a)) variants. On TSPLIB, DRAGON consistently outperforms ReEvo(c) across all size groups and also surpasses ReEvo(a) on all groups beyond 500, and achieving an average gap reduction of 0.86%. On CVRPLIB, where OR-Tools fails to find feasible solutions within time limit, DRAGON leads around 3.98% lower gaps than ReEvo(a) on all cases. Overall, DRAGON demonstrates the best average performance across routing benchmarks, validating its effectiveness and scalability.
@@ 4.3.2. Impact of LLM backbones
Figure 7. Comparison of LLM models by: (a) optimality gap, (b) number of API calls, (c) input tokens, (d) output tokens, (e) running time. We evaluate general-purpose models (gpt-4o, gpt-4.1) and reasoning-oriented models (o3, r1). Gray bars indicate incomplete results due to token limit violations.
@@ 4.4.1. Bin Packing Problem
Compared to routing, packing problems are generally easier for LLMs to reason about due to their more intuitive and straightforward constraints. For example, merging static segments ($\mathbf{s}$) into packed items requires less complex reasoning, allowing DRAGON to scale more effectively to larger instances. Results for EoH and EoH expert are reported directly from liu2024evolution, where running time was not provided. While DRAGON achieves the lowest optimality gap, it is slower than code-generation based methods. However, code-generation approaches come with high API costs and substantial time and resources to evolve or fine-tune high-quality heuristic functions. These overheads are not captured in current comparison. Each paradigm has its strengths. For large batches of similar instances, code-generation methods are effective as the learned heuristic can be reused across the dataset. In contrast, prompt-based approaches like DRAGON are better suited for dynamic environments where instance structures vary, offering flexibility without the need for retraining or code evolution.
@@ 5. Conclusion
This paper presents DRAGON, a novel framework that leverages LLM agents to solve large-scale COPs via an iterative decomposition reconstruction process inspired by divide-and-conquer. Experiments across diverse COP domains (e.g., routing and packing) show that DRAGON consistently achieves strong performance, especially on extreme large instances where traditional solvers or basic prompt-based methods often fail due to scalability or timeouts. By decomposing complex problems into context manageable subproblems, DRAGON enables LLMs to reason effectively over intricate structures, identifying promising regions, and reconstructing feasible global solutions. This training-free, modular, and solver-agnostic approach supports flexible constraint handling and adapts easily to diverse COPs without specialized algorithm design. Although not optimized for speed, DRAGON avoids expensive code search and offers a strong balance between solution quality, adaptability, and cost, making it particularly well-suited for dynamic and large-scale settings. Future work will focus on enhancing decomposition strategies, improving prompt robustness, and integrating hybrid solvers to explore promising regions more efficiently. Overall, DRAGON establishes a solid foundation toward general-purpose, interpretable, and scalable LLM-based optimization.
@@
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-017, AISG3-RP-2025-036-USNSF). For Supplementary Materials: https://arxiv.org/abs/2601.06502
@@ B.1. TSPLIB Results
We then present full results on benchmarking instances from TSPLIB in Tables A3A7. OPRO fails to produce valid solutions across all groups, as several hard instances exceed the runtime limit, leading to infinite gaps (inf). SGE consistently returns infeasible solutions (infe), likely due to the absence of explicit feasibility checks, as it was originally designed for small-scale problems (fewer than 30 nodes). LMEA can handle moderately larger instances, but suffers from high optimality gaps and fails on instances beyond 2k nodes, primarily due to prompt length constraints and generation timeouts.